Ken and Vince have $1002. Ken and Leon have $1342. The ratio of Vince to Leon's money is 2:7. How much does Ken have?
Answer by stanbon(75887)

Ken and Vince have $1002. Ken and Leon have $1342. The ratio of Vince to Leon's money is 2:7. How much does Ken have?
---
K + V = 1002
K + L = 1342
---
Subtract to get:
L - V = 342
=====================
V/L = 2/7
V = (2/7)L
==================
Substitute and solve for "L"
L - V = 342
L - (2/7)L = 342
(5/7)L = 342
L = 7*342/5
L = 478.8
---
Solve for "K":
K + L = 1342
K + 478.8 = 1342
Ken's amount = $863,20
===========================
Cheers,
Stan H.
==========